# Decentralized Federated Anomaly Detection in Smart Grids: A P2P Gossip Approach

### v1

@@ Abstract
The increasing security and privacy concerns in the Smart Grid sector have led to a significant demand for robust intrusion detection systems within critical smart grid infrastructure. To address the challenges posed by privacy preservation and decentralized power system zones with distinct data ownership, Federated Learning (FL) has emerged as a promising privacy-preserving solution which facilitates collaborative training of attack detection models without necessitating the sharing of raw data. However, FL presents several implementation limitations in the power system domain due to its heavy reliance on a centralized aggregator and the risks of privacy leakage during model update transmission. To overcome these technical bottlenecks, this paper introduces a novel decentralized federated anomaly detection scheme based on two main gossip protocols namely Random Walk and Epidemic. Our findings indicate that the Random Walk protocol exhibits superior performance compared to the Epidemic protocol, highlighting its efficacy in decentralized federated learning environments. Experimental validation of the proposed framework utilizing publicly available industrial control systems datasets demonstrates superior attack detection accuracy while safeguarding data confidentiality and mitigating the impact of communication latency and stragglers. Furthermore, our approach yields a notable 35% improvement in training time compared to conventional FL, underscoring the efficacy and robustness of our decentralized learning method.
Index Terms:
Anomaly Detection, Decentralized Federated Learning (DFL), Cyberattack, Internet of Things (IoT), Smart Grid
I Introduction
@@
The widespread adoption of Internet of Things (IoT) technology and advanced communication protocols within the Smart Grid (SG) paradigm has led to a significant surge in data volumes generated from edge devices. This vital power-related data play a crucial role in various key operations of the grid including demand management, grid monitoring and control, fault detection and isolation, etc. Consequently, the management of the overall grid infrastructure can be carried out more dynamically and efficiently. However, the increasing heterogeneity, diversity, and complexity within the power grid ecosystem present significant challenges to its integrity. Ensuring data integrity is now a priority for modern power systems due to the ubiquity of modern cyber-physical power grid components and advancements in information technology, which enable data processing and decision-making at the edge, as opposed to traditional cloud storage-based control centers.
@@
On another hand, whilst the digitisation of the modern-day power systems has mitigated the response issues to the ever-growing demand for energy, smart grids are still prone to several technical bottlenecks. Driven by the massive amounts of sensitive data generated, a wide range of the cyber-physical components of the SG have become potentially susceptible to zero-day cyber vulnerabilities by malicious actors, including but not limited to False Data Injection Attacks (FDIA), Time Synchronization Attacks, etc. Hence, it is preemptive to leverage state-of-the-art technological advances to develop effective countermeasures which defend SGs against such types of intrusions. Consequently, the use of Machine Learning and Deep Learning (DL) has emerged as driving technologies behind several breakthroughs in grid sabotage detection due to their ability to identify relevant complex patterns from diverse data sources.
@@
Despite their impressive contributions, such centralized anomaly detection schemes suffer from several drawbacks which include limited storage capabilities, communication network overhead and, foremost, privacy issues. Experts often disregard the fact that grid information contains sensitive data which carries potential privacy risks during transfer to centralized repositories for analytical tasks. Thus, to tackle such impediments, training deep learning models in a distributed fashion, essentially Federated Learning (FL) garnered momentum as a potential privacy-preserving alternative which enables edge-based on-device training without the need of sharing raw data thus mitigating the associated privacy risks. For instance, one of our previous works [1] in similar line proposed a computationally-efficient privacy-preserving federated scheme to differentiate between anomalous events and system disturbances. Other studies [2, 3, 4] have similarly leveraged FL techniques for misbehaviour detection within SG ecosystems. Such literature have explored the effectiveness of FL at detecting irregularities and deviations from normal operating conditions, thus, reinforcing the security and resilience of power grids.
@@
However, traditional FL-based intrusion detection schemes, whilst being focussed on its client-server-centric design principles as shown in Fig. 1A, exhibits certain drawbacks that impede its practical implementation within real-world power system scenarios. One such limitation is the blind reliance of client nodes on a centralized aggregator to coordinate the learning process and aggregate local model updates. This may eventually lead to potential privacy risks [5] and communication bottlenecks [1], especially in large-scale and geographically distributed environments. Furthermore, traditional FL approaches also lack behind in scalability, due to the centralized server becoming a single point of failure and expensive computational overhead with increasing number of client nodes. Such weaknesses motivate the exploration of Decentralized Federated Learning (DFL) whereby model aggregation and coordination occur both locally at the client nodes rather than requiring a central server, hence, reducing privacy concerns, improving scalability, and enhancing the robustness of distributed learning-based anomaly detection scenarios.
@@
In this work, we advocate for the effectiveness of a decentralized federated approach to misbehaviour detection in modern power systems in view of discriminating between malicious and natural power system irregularities, thereby providing enabling reliable operational support to energy grid operators. Specifically, the main contributions of this manuscript as follows: 1) Hybrid Anomaly Detection: We leverage the capabilities of state-of-the-art hybrid anomaly detection neural networks which combines the strength from both sequential modelling and representation learning, specifically a Transformer-Autoencoder (TAE) model to offer accurate and timely detection of intentionally crafted malicious events from natural power system disturbances. 2) Gossip-based Decentralized Federated Approach: To the best of our knowledge, we propose the first Random Walk gossip-based decentralized federated misbehavior detection approach which harnesses the collective intelligence of grid regions within a geographically decentralized power grid system. It also eliminates the need for a centralized aggregator, thereby mitigates the associated privacy concerns and scalability pitfalls associated with conventional federated anomaly detection methods. 3) Privacy-preserving Gradient Quantization: To further optimize the computational overhead of our proposed solution, we employ a differentially-private gradient quantizer, particularly DP-SIGNSGD, which significantly reduces the byte-size of the model updates propagated through the network whist improving its privacy aspects. 4) Lastly, we perform comprehensive empirical evaluations of our proposed frameworks on openly accessible Industrial Control Systems datasets to validate its performance computational overhead which yields a 94.2% accuracy and 35% cutback in training time against a conventional FL setup.
@@
The rest of this manuscript is structured as follows: In Section II, we briefly review the existing literature in relation to conventional FL and DFL. The proposed methodology, detection module and system model are thoroughly discussed in Section III. Simulation scenarios to evaluate the effectiveness of our proposed DFL scheme on publicly obtained databases are presented in Section IV. Finally, we conclude our article in Section V.
II Related Literature
@@
II-A Conventional FL-based Anomaly Detection & its Limitations
@@
Since its inception, FL has established itself as an emerging discipline that has achieved promising results in various SG-related applications whereby privacy is paramount. Taïk and Cherkaoui [6] introduced FL to the SG domain by training a LSTM model on a real-world Texas load consumption data to highlight the effectiveness of FL and its privacy-preserving guarantees in a demand load forecasting scenario. Another work in [7] proposed a horizontal federated load forecasting solution for Retail Energy Providers and achieved sufficient forecasting performance using the Solar Home Electricity Data from Ausgrid. Following its success, multiple studies including [8, 9, 10] continued in similar directions.
@@
Given its prominence and optimal results within demand forecasting, FL swiftly gained momentum in other facets of power systems, in particular, anomaly detection. For instance, researchers in [3] developed a federated temporal convolutional network for identifying power theft. Through comprehensive data-driven experiments using a real-energy consumption dataset, they asserted that their federated framework can attain high detection accuracy with reduced computational overhead. Similarly, Ashraf et al. [11] attained comparable performance for energy theft detection by validating federated conventional machine learning algorithms on real-world datasets. The works in [1, 12] proposed computationally-efficient differentially-private federated schemes to differentiate between anomalous events and system irregularities. Furthermore, authors in [13, 14, 15] developed FL-based frameworks for the detection of False Data Injection attacks in power systems with sufficient performance. More recently, to mitigate the effects of straggler client nodes and communication inefficiency in relation to conventional FL, the works in [16, 17, 18] proposed semi-asynchronous FL-based anomaly detection frameworks to discriminate between cyberattack events and natural power system disturbances which yielded improved attack detection performance and improved robustness versus stragglers.
@@
Whilst conventional FL has proved its effectiveness in the area of anomaly detection for power systems, it comes with inherent limitations that necessitate further investigations. One such major drawback is its heavy reliance on a centralized aggregator which introduces potential privacy and security hazards during the transmission of sensitive model updates to and from client nodes. Furthermore, the communication overhead associated with the bi-directional transmission of model gradients from numerous edge devices to the central aggregator can be computationally expensive, thus leading to scalability issues, particularly in scenarios with a large number of client nodes or when operating in resource-constrained environments. Hence, it is imperative to devise a solution to address these technical bottlenecks.
II-B Decentralized Federated Anomaly Detection
@@
In the context of power system scenarios, the novelty of Decentralized Federated Learning (DFL) emerges as a vital solution to overcome the aforementioned technical challenges associated with conventional FL-based anomaly detection. DFL is a recently surfaced machine learning paradigm which combines the principles of FL with a decentralized architecture. In conventional FL, a central aggregator orchestrates the model training process by aggregating model updates from multiple client nodes. In contrast, model training is distributed across multiple client nodes during DFL. Each node independently trains a local model using its own data and periodically exchanges model updates or gradients with other client nodes within the network. In the context of power systems, Giuseppi et al. [19] devised a decentralized version of FederatedAveraging (DecFedAvg) algorithm for non-intrusive load monitoring in smart energy communities and validated the effectiveness of their innovative approach using REFIT dataset which yielded low Mean Absolute Error (MAE) values. Another work by Gao et al. [20] presented a DFL-based residential building load forecasting framework with a gradient selection mechanism and achieved a 97% accuracy on Pecan Street Dataset. Remarkably, only a limited number of studies have explored the use of DFL within the smart grid domain with a notable absence of studies specifically addressing the issue of anomaly detection. This evident gap in existing literature highlights the need for additional investigation, prompting our motivation. Therefore, this manuscript aims to address this gap by proposing a novel approach to anomaly detection in power systems by decentralizing the FL process using a gossip-based approach to eliminate the dependency on a centralized aggregator.
III Proposed Method
@@
III-A Power System Model & Assumptions
@@
Throughout this manuscript, we consider a geographically sparse and decentralized electrical grid system that is segmented into $N$ number of power grid regions where $N\in\mathbb{R}^{+}$. As depicted in Fig 2, we briefly introduce the primary actors of power system model as in the following: 1) SCADA Sub-systems: The SCADA sub-systems are the edge-based client nodes equipped with sophisticated data acquisition technological capabilities which play a very crucial role in gathering and overseeing the electrical data sensed by each IoT-based device (including relays, Remote Terminal Units (RTUs), etc.) within a power grid region. Furthermore, the SCADA sub-systems will function as the DFL nodes whereby they will each be training a model locally using their own data repository and update the one-bit quantized model updates to selected peers within the network. Lastly, the SCADA sub-systems are responsible for the aggregation of the received updates from their peers with their own local model parameters. For this given scenario, we assume that the SCADA sub-systems are honest-but-curious, i.e, they strictly adhere to the protocols of DFL without exhibiting malicious behaviours to ensure data integrity and privacy, model quality, fairness, etc. 2) Grid Regions: The power grid regions are equipped with a variety of decentralized monitoring sensoring devices including relays, synchrophasors, RTUs, etc. and are responsible for the continuous sensing of grid-related data such as current phase angle, voltage magnitude, etc. The Intelligent Electronic Devices (IEDs) are presumably connected to the SCADA sub-systems via high speed data transmission networks built upon DNP3/IEC 61850 protocols. Furthermore, we assume that the grid regions are interconnected for operational purposes using wireless mesh networking protocols including Zigbee, etc.
III-B Proposed Anomaly Detection Mechanism
@@
As previously mentioned in Section II above, previous works in similar field have employed a number of anomaly detection techniques including statistical approaches, traditional ML algorithms, DL neural networks, etc. However, in this manuscript, we formulate the problem of cyber attack detection using a Transformer-Autoencoder (TAE) model as illustrated in Fig. 3. TAEs are powerful hybrid neural networks that combines the elements of both transformer and autoencoder models, leveraging the strengths of sequence modelling and representation learning, with the objective of achieving identity mapping between inputs and outputs, whilst capturing complex patterns and dependencies from the data [21]. Autoencoders learn representations in data by passing an input data vector $v_{i}$ through an encoder $f_{\alpha}$ which maps each input $v_{i}$ to a latent space representation $r^{(j)}$ which is essentially a low-dimensional and compressed knowledge representation of the original input such that $r^{(j)}=f_{\alpha}(v_{i})=a(Wv_{i}+b)$ where $a$ is an element-wise activation function, $W$ is the weight matrix and $b$ is the bias vector of the encoder. On the other hand, the decoder $g_{\alpha}$ is responsible for the reconstruction of the original input $v_{i}$, with minimal reconstruction error, from $r^{(j)}$ through $v_{j}=g_{\alpha}(v_{i})=a(W^{\prime}v_{i}+b^{\prime})$ where $v_{j}$ is the reconstructed input (output of decoder). During the federated training process, the model parameters $W$, $W^{\prime}$, $b$ and $b^{\prime}$ are iteratively improved to optimize the reconstruction error $\ell(v_{i},v_{j})$ such that $\alpha,\alpha^{\prime}=argmin_{\alpha,\alpha^{\prime}}\ell(v_{i},v_{j})$.
@@
III-C Gradient Quantization
During federated training, the canonical objective of the TAE model is minimize the objective function $argmin_{\alpha,\alpha^{\prime}}\ell(v_{i},v_{j})$. Typically, the search for the optimal parameter values of this objective function within a FL/DFL set-up can be accomplished by employing several optimization algorithms including Stochastic Gradient Descent (SGD), Adam, etc. However, due to the requirement of bi-directional communication amongst grid regions, the size and frequency of the local gradient updates increases which hinders the scalability of our proposed approach to increasing number of grid regions. Additionally, the sharing of raw model updates poses significant privacy risks to such a critical system, this non-conformant to confidentiality guarantees of FL [5].
Therefore, to overcome such technical bottlenecks, we propose to employ a gradient quantization scheme, specifically DP-SIGNSGD [1], which approximates the TAE model gradients into low-precision values to reduce communication bandwidth. After aggregation with previous model updates, each decentralized grid region $k$ executes a zero-mean normalization of its computed gradient $g_{k}^{t}$ such that $\overline{g}_{k}^{t}=g_{k}^{t}-\mu_{k}^{t}$ for uniform distribution, followed by a compression of the normalized gradients by each grid region $k$ through a one-bit differentially-private quantizer which eventually captures the sign of $\overline{g}_{k}^{t}$, such that $\hat{g}_{k}^{t}=dpsign(\overline{g}_{k}^{t})$. The quantized gradient $\hat{g}_{k}^{t}$ from one grid region is then sent to a randomly selected neighbour node. The latest model gradient is achieved by updating the new weight parameters such that $W^{t+1}=W^{t}-\eta^{t}\hat{g}$ where $\eta^{t}\in(0,1)$ represents the learning rate. This gradient quantization technique, as detailed in Algorithm 1, significantly decreases the byte size of the gradient updates exchanged during the gossip-based model training.
III-D Gossip-based DFL Framework

### v2

@@ Abstract
Amidst escalating concerns regarding security and privacy within the Smart Grid domain, the need for robust intrusion detection mechanisms in critical energy infrastructure has surged in recent times. To address the challenges posed by privacy preservation and decentralized power zones with distinct data ownership, Federated Learning (FL) has emerged as a promising privacy-preserving solution which facilitates collaborative training of attack detection models without necessitating the sharing of raw data. However, FL presents several implementation limitations in the power system domain due to its heavy reliance on a centralized aggregator and the risks of privacy leakage during model update transmission. In response to the technical bottlenecks, this paper introduces a novel decentralized federated anomaly detection scheme based on two main gossip protocols namely Random Walk and Epidemic. Our findings indicate that the Random Walk protocol exhibits superior performance compared to the Epidemic protocol, highlighting its efficacy in decentralized federated learning environments. Experimental validation of the proposed framework utilizing publicly available industrial control systems datasets demonstrates superior attack detection accuracy while safeguarding data confidentiality and mitigating the impact of communication latency and stragglers. Moreover, a notable 35% improvement in training time against conventional FL highlights the efficacy and robustness of our decentralized learning approach.
keywords:
[…]
1 Introduction
@@
The large-scale prevalence of Internet of Things (IoT) technology and cutting-edge communication protocols within the Smart Grid (SG) paradigm has provoked a significant surge in data volumes emanating from edge devices. Such vital power-related data play a crucial role in various key operations of the grid including demand management, grid monitoring and control, fault detection and isolation, etc. Consequently, the management of the overall grid infrastructure can be carried out more dynamically and efficiently. However, the increasing heterogeneity, diversity, and complexity within the power grid ecosystem present significant challenges to its integrity. In modern times, ensuring data integrity is a priority for modern power systems which is due to the ubiquity of modern cyber-physical power grid components alongside advancements in information technology, which now enable data processing and decision-making at the edge, in contrast to traditional cloud storage-based control centers.
@@
On another hand, whilst the digitisation of the modern-day power systems has managed to alleviate the response issues to the insatiable demand for energy, smart grids are still prone to several technical bottlenecks. Empowered by the massive amounts of sensitive data generated, a wide range of the cyber-physical components of the SG has become potential zero-day vulnerability susceptible to cyber threats by malicious actors, including but not limited to False Data Injection Attacks (FDIA), Time Synchronization Attacks, etc. Hence, it is preemptive to leverage state-of-the-art technological advances to develop effective countermeasures to defend SGs against such types of intrusions. Consequently, the use of Machine Learning and Deep Learning (DL) emerged as driving forces behind several breakthroughs in grid sabotage detection due to their prowess at identifying relevant complex patterns from diverse data sources.
@@
Despite their remarkable contributions, such focalized anomaly detection schemes suffer from several shortcomings which include limited storage capabilities, communication network overhead and, foremost, privacy issues. Experts frequently disregard the fact that grid information encompasses sensitive data which carries potential privacy risks during transfer to centralized repositories for analytical tasks. Thus, to tackle such impediments, training deep learning models in a distributed fashion, essentially Federated Learning (FL) garnered momentum as a plausible privacy-preserving alternative which enables edge-based on-device training bereft of sharing raw data thus mitigating the associated privacy risks. For instance, one of our previous works Husnoo et al. (2023d) in similar line proposed a computationally-efficient privacy-preserving federated scheme to differentiate between anomalous events and system disturbances. Other studies Su et al. (2022); Wen et al. (2022); Huong et al. (2021) have similarly leveraged FL techniques for misbehaviour detection within SG ecosystems. Such literature have explored the effectiveness of FL at detecting irregularities and deviations from normal operating conditions, thus, reinforcing the security and resilience of power grids.
@@
However, traditional FL-based intrusion detection schemes, whilst being focussed on its client-server-centric design principles as shown in Fig. 1A, exhibits certain drawbacks that impede its practical implementation within real-world power system scenarios. One such limitation is the blind reliance of client nodes on a centralized aggregator to coordinate the learning process and aggregate local model updates. This may eventually lead to potential privacy risks Husnoo et al. (2024) and communication bottlenecks Husnoo et al. (2023d), especially in large-scale and geographically distributed environments. Furthermore, traditional FL approaches also lack behind in scalability, due to the centralized server becoming a single point of failure and expensive computational overhead with increasing number of client nodes. Such weaknesses motivate the exploration of Decentralized Federated Learning (DFL) whereby model aggregation and coordination occur both locally at the client nodes rather than requiring a central server, hence, reducing privacy concerns, improving scalability, and enhancing the robustness of distributed learning-based anomaly detection scenarios.
@@
In this work, we advocate for the effectiveness of a decentralized federated approach to misbehaviour detection in modern power systems in view of discriminating between malicious and natural power system irregularities, thereby providing enabling reliable operational support to energy grid operators. Specifically, the main contributions of this manuscript are laid out in the following: 1) We leverage the capabilities of state-of-the-art hybrid anomaly detection neural networks which combines the strength from both sequential modelling and representation learning, specifically a Transformer-Autoencoder (TAE) model to offer accurate and timely detection of intentionally crafted malicious events from natural power system disturbances. 2) To the best of our knowledge, we propose the first gossip-based decentralized federated misbehavior detection approach which harnesses the collective intelligence of grid regions within a geographically decentralized power grid system while eliminating the need for a centralized aggregator, thereby mitigating the associated privacy concerns and scalability pitfalls associated with conventional federated anomaly detection methods. 3) To further optimize the computational overhead of our proposed solution, we employ a differentially-private gradient quantizer, particularly DP-SIGNSGD, which significantly reduces the byte-size of the model updates propagated through the network whist improving its privacy aspects. 4) Lastly, we perform comprehensive empirical evaluations of our proposed frameworks on openly accessible Industrial Control Systems dataset to validate its performance computational overhead which yields a 94.2% accuracy and 35% cutback in training time against a conventional FL setup.
@@
The rest of this manuscript is structured as follows: In Section 2, we briefly review the existing literature in relation to conventional FL and DFL. The proposed methodology, detection module and system model are thoroughly discussed in Section 3. Simulation scenarios to evaluate the effectiveness of our proposed DFL scheme on publicly obtained databases are presented in Section 4. Finally, we conclude our article in Section 5.
2 Related Literature
@@
2.1 Conventional FL-based Anomaly Detection & its Limitations
@@
Since its inception, FL has positioned itself as an emerging discipline that has achieved promising results in various SG-related applications whereby privacy is paramount. Taïk and Cherkaoui Taïk and Cherkaoui (2020) introduced FL to the SG domain by training a LSTM model on a real-world Texas load consumption data to highlight the effectiveness of FL and its privacy-preserving guarantees in a demand load forecasting scenario. Another work in Husnoo et al. (2022) proposed a horizontal federated load forecasting solution for Retail Energy Providers and achieved sufficient forecasting performance using the Solar Home Electricity Data from Ausgrid. Following its success, multiple studies including Fekri et al. (2022); Briggs et al. (2022); Gholizadeh and Musilek (2022) continued in similar directions.
@@
Given its prominence and optimal results within demand forecasting, FL swiftly gained momentum in other facets of power systems, in particular, anomaly detection. For instance, researchers in Wen et al. (2022) developed a federated temporal convolutional network for identifying power theft. Through comprehensive data-driven experiments using a real-energy consumption dataset, they asserted that their federated framework can attain high detection accuracy with reduced computational overhead. Similarly, Ashraf et al. Ashraf et al. (2022) attained comparable performance for energy theft detection by validating federated conventional machine learning algorithms on real-world datasets. The works in Husnoo et al. (2023d, a) proposed computationally-efficient differentially-private federated schemes to differentiate between anomalous events and system irregularities. Furthermore, authors in Zhao et al. (2022); Lin et al. (2022); Li et al. (2022) developed FL-based frameworks for the detection of False Data Injection attacks in power systems with sufficient performance. More recently, to mitigate the effects of straggler client nodes and communication inefficiency in relation to conventional FL, the works in Husnoo et al. (2023e, c, b) proposed semi-asynchronous FL-based anomaly detection frameworks to discriminate between cyberattack events and natural power system disturbances which yielded improved attack detection performance and improved robustness versus stragglers.
@@
Whilst conventional FL has proved its effectiveness in the area of anomaly detection for power systems, it is accompanied with inherent limitations that necessitate further investigations. One such major drawback is its heavy reliance on a centralized aggregator which introduces potential privacy and security hazards during the transmission of sensitive model updates to and from client nodes. Furthermore, the communication overhead associated with the bi-directional transmission of model gradients from numerous edge devices to the central aggregator can be computationally expensive, thus leading to scalability issues, particularly in scenarios with a large number of client nodes or when operating in resource-constrained environments. Hence, it is imperative to devise a solution to address these technical bottlenecks.
2.2 Decentralized Federated Anomaly Detection
@@
In the context of power system scenarios, the novelty of Decentralized Federated Learning (DFL) emerges as a vital solution to overcome the aforementioned technical challenges associated with conventional FL-based anomaly detection. DFL is a recently surfaced machine learning paradigm which combines the principles of FL with a decentralized architecture. In conventional FL, a central aggregator orchestrates the model training process by aggregating model updates from multiple client nodes. In contrast, model training is distributed across multiple client nodes during DFL. Each node independently trains a local model using its own data and periodically exchanges model updates or gradients with other client nodes within the network. In the context of power systems, Giuseppi et al. Giuseppi et al. (2022) devised a decentralized version of FederatedAveraging (DecFedAvg) algorithm for non-intrusive load monitoring in smart energy communities and validated the effectiveness of their innovative approach using REFIT dataset which yielded low Mean Absolute Error (MAE) values. Another work by Gao et al. Gao et al. (2021) presented a DFL-based residential building load forecasting framework with a gradient selection mechanism and achieved a 97% accuracy on Pecan Street Dataset. Remarkably, only a limited number of studies have explored the use of DFL within the smart grid domain with a notable absence of studies specifically addressing the issue of anomaly detection. This evident gap in existing literature highlights the need for additional investigation, prompting our motivation. Therefore, this manuscript aims to address this gap by proposing a novel approach to anomaly detection in power systems by decentralizing the FL process using a gossip-based approach to eliminate the dependency on a centralized aggregator.
3 Proposed Method
@@
3.1 Power System Model & Assumptions
@@
Throughout this manuscript, we consider a geographically decentralized power grid system that is segmented into $N$ number of grid regions where $N\in\mathbb{R}^{+}$. As depicted in Fig 2, we briefly introduce the primary actors of power system model as follows: 1) SCADA Sub-systems: The SCADA sub-systems are the edge-based client nodes equipped with sophisticated data acquisition technologies which play a crucial role in gathering and overseeing the data sensed by each sensing device (including relays, Remote Terminal Units (RTUs), etc.) within a power grid region. Furthermore, the SCADA sub-systems will function as DFL nodes whereby they will each train a model locally using their own data repository and update the compressed gradients to selected peers within the network. Lastly, the SCADA sub-systems are responsible for the aggregation of the received updates from their peers with their own local model parameters. For this given scenario, we assume that the SCADA sub-systems are honest-but-curious, i.e, they strictly adhere to the protocols of DFL without exhibiting malicious behaviours to ensure data integrity and privacy, model quality, fairness, etc. 2) Grid Regions: The power grid regions are equipped with a variety of decentralized monitoring sensoring devices including relays, synchrophasors, RTUs, etc. and are responsible for the continual sensing of grid-related data such as current phase angle, voltage magnitude, etc. The Intelligent Electronic Devices (IEDs) are presumably connected to the SCADA sub-systems via high speed data transmission networks built upon DNP3/IEC 61850 protocols. Furthermore, we assume that the grid regions are interconnected for operational purposes using wireless mesh networking protocols including Zigbee, etc.
3.2 Proposed Anomaly Detection Mechanism
@@
As previously mentioned in Section 2 above, previous works in similar field have employed a number of anomaly detection techniques including statistical approaches, traditional ML algorithms, DL neural networks, etc. However, in this manuscript, we formulate the problem of cyber attack detection using a Transformer-Autoencoder (TAE) model as illustrated in Fig. 3. TAEs are powerful hybrid neural networks that combines the elements of both transformer and autoencoder models, leveraging the strengths of sequence modelling and representation learning, with the objective of achieving identity mapping between inputs and outputs, whilst capturing complex patterns and dependencies from the data Shabanpour et al. (2023). Autoencoders learn representations in data by passing an input data vector $v_{i}$ through an encoder $f_{\alpha}$ which maps each input $v_{i}$ to a latent space representation $r^{(j)}$ which is essentially a low-dimensional and compressed knowledge representation of the original input such that $r^{(j)}=f_{\alpha}(v_{i})=a(Wv_{i}+b)$ where $a$ is an element-wise activation function, $W$ is the weight matrix and $b$ is the bias vector of the encoder. On the other hand, the decoder $g_{\alpha}$ is responsible for the reconstruction of the original input $v_{i}$, with minimal reconstruction error, from $r^{(j)}$ through $v_{j}=g_{\alpha}(v_{i})=a(W^{\prime}v_{i}+b^{\prime})$ where $v_{j}$ is the reconstructed input (output of decoder). During the federated training process, the model parameters $W$, $W^{\prime}$, $b$ and $b^{\prime}$ are iteratively improved to optimize the reconstruction error $\ell(v_{i},v_{j})$ such that $\alpha,\alpha^{\prime}=argmin_{\alpha,\alpha^{\prime}}\ell(v_{i},v_{j})$.
@@
3.3 Gradient Quantization
During federated training, the canonical objective of the TAE model is minimize the objective function $argmin_{\alpha,\alpha^{\prime}}\ell(v_{i},v_{j})$. Typically, the search for the optimal parameter values of this objective function within a FL/DFL set-up can be accomplished by employing several optimization algorithms including Stochastic Gradient Descent (SGD), Adam, etc. However, due to the requirement of bi-directional communication amongst grid regions, the size and frequency of the local gradient updates increases which hinders the scalability of our proposed approach to increasing number of grid regions. Additionally, the sharing of raw model updates poses significant privacy risks to such a critical system, this non-conformant to confidentiality guarantees of FL Husnoo et al. (2024).
Therefore, to overcome such technical bottlenecks, we propose to employ a gradient quantization scheme, specifically DP-SIGNSGD Husnoo et al. (2023d), which approximates the TAE model gradients into low-precision values to reduce communication bandwidth. After aggregation with previous model updates, each decentralized grid region $k$ executes a zero-mean normalization of its computed gradient $g_{k}^{t}$ such that $\overline{g}_{k}^{t}=g_{k}^{t}-\mu_{k}^{t}$ for uniform distribution, followed by a compression of the normalized gradients by each grid region $k$ through a one-bit differentially-private quantizer which eventually captures the sign of $\overline{g}_{k}^{t}$, such that $\hat{g}_{k}^{t}=dpsign(\overline{g}_{k}^{t})$. The quantized gradient $\hat{g}_{k}^{t}$ from one grid region is then sent to a randomly selected neighbour node. The latest model gradient is achieved by updating the new weight parameters such that $W^{t+1}=W^{t}-\eta^{t}\hat{g}$ where $\eta^{t}\in(0,1)$ represents the learning rate. This gradient quantization technique, as detailed in Algorithm 1, significantly decreases the byte size of the gradient updates exchanged during the gossip-based model training.
3.4 Gossip-based DFL Framework
To enable federated discrimination of attack events from non-anomalous ones, we propose to implement two types of gossip protocols: 1) Random Gossip randomly selects client nodes within a network for communication. During each communication round, each $k$ computes its local gradient $\nabla f_{k}(Z^{t})$ using its training dataset $D_{k}$ and aggregates this gradient with that received from a randomly chosen peer node. This aggregated gradient is then normalized and quantized before being transmitted to another randomly selected peer node. This process repeats iteratively, facilitating decentralized communication and collaboration among nodes. 2) Epidemic Gossip, on the other hand, operates on the principle of spreading information to all client nodes within a network. Each $k$ communicates with a fixed subset of its neighboring $k$, known as its Epidemic Radius. Similar to random gossip, nodes compute, aggregate, normalize, and quantize gradients before transmitting them to neighboring nodes. However, unlike random gossip, Epidemic gossip ensures that each node eventually receives the information from all other nodes in the network. By leveraging random gossip, nodes can efficiently exchange information without requiring explicit knowledge of the network topology. This enhances scalability and adaptability to changes in network conditions. On the other hand, Epidemic gossip ensures that all nodes eventually receive the information, thereby promoting information dissemination and consensus among nodes.